# NovoBench: Benchmarking Deep Learning-based De Novo Peptide Sequencing Methods in Proteomics

### v1

@@ Abstract
Tandem mass spectrometry has played a pivotal role in advancing proteomics, enabling the high-throughput analysis of protein composition in biological tissues. Many deep learning methods have been developed for de novo peptide sequencing task, i.e., predicting the peptide sequence for the observed mass spectrum. However, two key challenges seriously hinder the further advancement of this important task. Firstly, since there is no consensus for the evaluation datasets, the empirical results in different research papers are often not comparable, leading to unfair comparison. Secondly, the current methods are usually limited to amino acid-level or peptide-level precision and recall metrics. In this work, we present the first unified benchmark NovoBench for de novo peptide sequencing, which comprises diverse mass spectrum data, integrated models, and comprehensive evaluation metrics. Recent impressive methods, including DeepNovo, PointNovo, Casanovo, InstaNovo, AdaNovo and $\pi$-HelixNovo are integrated into our framework. In addition to amino acid-level and peptide-level precision and recall, we evaluate the models’ performance in terms of identifying post-tranlational modifications (PTMs), efficiency and robustness to peptide length, noise peaks and missing fragment ratio, which are important influencing factors while seldom be considered. Leveraging this benchmark, we conduct a large-scale study of current methods, report many insightful findings that open up new possibilities for future development. The benchmark will be open-sourced to facilitate future research and application.
@@ 1 Introduction
Proteomics, the study of proteins within biological systems, relies heavily on mass spectrometry for protein identification aebersold2003mass . Traditional methods use existing databases to match observed peptide fragments with known sequences. However, these methods may miss novel or modified peptides not present in the databases ma2003peaks ; ma2015novor ; nesvizhskii2010survey . De novo peptide sequencing offers a solution by directly annotating mass spectra to reconstruct peptide sequences without relying on databases. By circumventing the need for predefined databases, de novo sequencing enables researchers to uncover new peptides and investigate post-translational modifications (PTMs), contributing to a deeper understanding of cellular processes uzozie2018advancing and disease mechanisms mann2003proteomic . Here, PTMs refers to chemically modified version of 20 naturally-occurring amino acids, influencing many critical biomolecular processes including central enzyme activities, protein turnover, and DNA repair ramazi2021post ; duan2015roles ; pejaver2014structural . Deep learning has been widely used in the de novo peptide sequencing, where researchers “translate” the observed mass spectrum to peptide sequence using the encoder-decoder architectures tran2017novo ; qiao2021computationally ; yilmaz2022novo ; mao2023mitigating ; xia2024adanovo ; liu2023accurate .
@@ 6 Influencing Factors
In this section, we investigate the influence of various factors on performance of de novo peptide sequence models to evaluate the robustness of all models under varying degrees of missing fragmentation, noise peaks and peptide lengths.
@@ 7.1 Experimental Settings
The model hyperparameters used in this paper are consistent with those in the original papers. We set the batch size to 32 and train the models for 30 epochs. DeepNovo and PointNovo were validated every 3000 steps, while the other models were validated every 50000 steps. The model with the lowest validation loss was selected for testing. All experimental results were obtained using a Nvidia A100 GPU (80GB). Please refer to the supplementary material for the detailed settings of each model.
@@ 7.2 Main Results
De novo peptide sequencing. As can be observed from Table 2, de novo sequencing models reveals varied performance at both the amino acid and peptide levels across different datasets. At the amino acid level, $\pi$-HelixNovo stands out with the high precision and recall across the Nine-species and HC-PT datasets, achieving precision and recall values of 0.765 and 0.758, and 0.588 and 0.582, respectively. DeepNovo exhibits the highest precision (0.492) on the Seven-species dataset but shows less impressive performance on the other datasets. PointNovo demonstrated the highest precision and recall (0.623 and 0.622) on the HC-PT dataset. However, compared to its performance on the high-resolution datasets Nine-species and HC-PT, the model did not perform as well on the low-resolution Seven-species dataset. CasaNovo shows consistent but average performance across all datasets, whereas InstaNovo demonstrates the weakest performance with notably low precision and recall values in all cases. Notably, we found that InstaNovo have a lower training loss than CasaNovo, which also uses a Transformer architecture, across all three datasets. However, its performance on the validation set was not as good, limiting its generalization capability. AdaNovo performs well on the Nine-species dataset, with a recall of 0.709, but is otherwise unremarkable.
@@
At the peptide level, $\pi$-HelixNovo again demonstrates superior performance, particularly in the Seven-species, with the highest precision and AUC values. In contrast, DeepNovo and CasaNovo show moderate performance, with DeepNovo slightly edging out in precision for the Seven-species dataset (0.204). PointNovo excels on the HC-PT dataset, achieving the highest precision (0.419) and AUC (0.373) among all models. InstaNovo remains the weakest model at the peptide level, with significantly lower precision and AUC values compared to other models. AdaNovo performs best on the nine-species dataset, achieving the highest AUC (0.469). Regarding confidence score, as shown in Table 4, $\pi$-HelixNovo has the highest confidence score, while PointNovo has the lowest. The confidence scores of Transformer-based models are notably higher compared to traditional methods.
@@
Identifying PTMs. The results in Table 3 show that DeepNovo achieves the highest PTM recall for the Seven-species dataset (0.373) and performs well on both the Nine-species and HC-PT datasets. PointNovo excels in the HC-PT dataset, attaining the highest PTM recall (0.740) and precision (0.676). CasaNovo demonstrates high PTM precision on the Nine-species dataset (0.706). AdaNovo performs competitively, achieving the second-highest PTM recall for the nine-species dataset (0.570) and demonstrating high precision across the datasets, particularly with a PTM precision of 0.448 for the seven-species dataset. $\pi$-HelixNovo exhibits consistent performance across all datasets, securing the top PTM recall for the Nine-species dataset (0.598) and achieving notable precision scores.
@@
Overall, $\pi$-HelixNovo consistently performs well across all datasets and metrics, demonstrating the effectiveness of complementary spectrum in de novo peptide sequencing task. Moreover, comparing results between CasaNovo and InstaNovo highlights the ineffectiveness of integrating precursor information into the decoder cross attention. Additionly, traditional deep learning-based methods remain highly competitive, and designing a better model architecture is still worth exploring.
@@
Table 3: Empirical comparison of de novo sequencing models in terms of identifying PTMs. The best results and the second best are highlighted with bold and underline, respectively.
@@
Table 4: Confidence score of different models on various datasets.The best results and the second best are highlighted with bold and underline, respectively.

### v2

@@ Abstract
Tandem mass spectrometry has played a pivotal role in advancing proteomics, enabling the high-throughput analysis of protein composition in biological tissues. Many deep learning methods have been developed for de novo peptide sequencing task, i.e., predicting the peptide sequence for the observed mass spectrum. However, two key challenges seriously hinder the further advancement of this important task. Firstly, since there is no consensus for the evaluation datasets, the empirical results in different research papers are often not comparable, leading to unfair comparison. Secondly, the current methods are usually limited to amino acid-level or peptide-level precision and recall metrics. In this work, we present the first unified benchmark NovoBench for de novo peptide sequencing, which comprises diverse mass spectrum data, integrated models, and comprehensive evaluation metrics. Recent impressive methods, including DeepNovo, PointNovo, Casanovo, InstaNovo, AdaNovo and $\pi$-HelixNovo are integrated into our framework. In addition to amino acid-level and peptide-level precision and recall, we evaluate the models’ performance in terms of identifying post-tranlational modifications (PTMs), efficiency and robustness to peptide length, noise peaks and missing fragment ratio, which are important influencing factors while seldom be considered. Leveraging this benchmark, we conduct a large-scale study of current methods, report many insightful findings that open up new possibilities for future development. The code is available at https://github.com/Westlake-OmicsAI/NovoBench.
@@ 1 Introduction
Proteomics, the study of proteins within biological systems, relies heavily on mass spectrometry for protein identification aebersold2003mass . Traditional methods use existing databases to match observed peptide fragments with known sequences. However, these methods may miss novel or modified peptides not present in the databases ma2003peaks ; ma2015novor ; nesvizhskii2010survey . De novo peptide sequencing offers a solution by directly annotating mass spectra to reconstruct peptide sequences without relying on databases. By circumventing the need for predefined databases, de novo sequencing enables researchers to uncover new peptides and investigate post-translational modifications (PTMs), contributing to a deeper understanding of cellular processes uzozie2018advancing and disease mechanisms mann2003proteomic . Here, PTMs refers to chemically modified version of 20 naturally-occurring amino acids, influencing many critical biomolecular processes including central enzyme activities, protein turnover, and DNA repair ramazi2021post ; duan2015roles ; pejaver2014structural . Deep learning has been widely used in the de novo peptide sequencing, where researchers “translate” the observed mass spectrum to peptide sequence using the encoder-decoder architectures tran2017novo ; qiao2021computationally ; yilmaz2022novo ; mao2023mitigating ; xia2024adanovo ; liu2023accurate ; xia2024bridging .
@@ 6 Influencing Factors
In this section, we investigate the influence of various factors on performance of de novo peptide sequence models to evaluate the robustness of all models against varying degrees of missing fragmentation, noise peaks and peptide lengths.
@@ 7.1 Experimental Settings
The model hyperparameters used in this paper are consistent with those in the original papers. We set the batch size to 32 and train the models for 30 epochs. DeepNovo and PointNovo were validated every 3000 steps, while the other models were validated every 50000 steps. The model with the lowest validation loss was selected for testing. We used the official pre-trained Instanovo model for testing on the Nine-species and HC-PT datasets, while the remaining models were trained from scratch on three datasets. All experimental results were obtained using a Nvidia A100 GPU (80GB). Please refer to the supplementary material for the detailed settings of each model.
@@ 7.2 Main Results
De novo peptide sequencing. As can be observed from Table 2, de novo sequencing models reveals varied performance at both the amino acid and peptide levels across different datasets. At the amino acid level, $\pi$-HelixNovo stands out with the high precision and recall across the Seven-species and Nine-species datasets, achieving precision and recall values of 0.481 and 0.472, and 0.765 and 0.758, respectively. InstaNovo achieved the highest precision (0.714) and recall (0.712) on the HC-PT dataset. DeepNovo exhibits the highest precision (0.492) on the Seven-species dataset but shows less impressive performance on the other datasets. PointNovo demonstrated the second-highest precision and recall (0.623 and 0.622) on the HC-PT dataset. However, compared to its performance on the high-resolution datasets Nine-species and HC-PT, the model did not perform as well on the low-resolution Seven-species dataset. CasaNovo shows consistent but average performance across all datasets. AdaNovo performs well on the Nine-species dataset, with a recall of 0.709, but is otherwise unremarkable.
@@
At the peptide level, InstaNovo demonstrates superior performance on the Nine-species and HC-PT datasets, with the highest precision and AUC values. $\pi$-HelixNovo achieved the best precision and AUC value on the Seven-species dataset. In contrast, DeepNovo and CasaNovo show moderate performance, with DeepNovo slightly edging out in precision for the Seven-species dataset (0.204). PointNovo excels on the HC-PT dataset, achieving the second-highest precision (0.419) and AUC (0.373) among all models. AdaNovo achieves the second-highest AUC (0.469) on Nine-species dataset. Regarding confidence score, as shown in Table 4, $\pi$-HelixNovo has the highest confidence score, while InstaNovo has the lowest. Additionally,the confidence scores of Transformer-based models are notably higher compared to traditional methods.
@@
Identifying PTMs. The results in Table 3 show that DeepNovo achieves the highest PTM recall for the Seven-species dataset (0.373) and performs well on both the Nine-species and HC-PT datasets. PointNovo excels in the HC-PT dataset, attaining the second-highest PTM recall (0.740) and precision (0.676). CasaNovo demonstrates high PTM precision on the Nine-species dataset (0.706). InstaNovo achieves the highest performance on HC-PT dataset. AdaNovo performs competitively, achieving the second-highest PTM recall for the nine-species dataset (0.570) and a PTM precision of 0.448 for the Seven-species dataset. $\pi$-HelixNovo exhibits consistent performance across all datasets, securing the top PTM recall for the Nine-species dataset (0.598) and achieving notable precision scores.
@@
Overall, $\pi$-HelixNovo consistently performs well across all datasets and metrics, demonstrating the effectiveness of complementary spectrum in de novo peptide sequencing task. InstaNovo demonstrates strong performance on both the Nine-species and HC-PT datasets, highlighting its robust capabilities after pre-training on large-scale, high-resolution mass spectrometry data. Additionally, traditional deep learning-based methods remain highly competitive, and designing a better model architecture is still worth exploring.
@@
Table 3: Empirical comparison of de novo sequencing models in terms of identifying PTMs. The best and the second best are highlighted with bold and underline, respectively.
@@
Table 4: Confidence score of different models on various datasets. The symbol * identifies the pre-trained model. The best and the second best are highlighted with bold and underline, respectively.
@@ 8 Conclusion
9 Limitation and Future Work
Given that Data-Independent Acquisition (DIA) data differs from Data-Dependent Acquisition (DDA) data and that there are currently fewer models ebrahimi2024transformerbasednovopeptidesequencing ; Tran2018DeepLE available for DIA, the availability of open-source DIA models is limited. Furthermore, de novo peptide sequencing models have already been widely applied to DDA mass spectrometry data, so our work is currently limited to DDA data. In the future, we will support DIA mass spectrometry data and DIA-specific models.
10 Acknowledgements
This work was supported by National Science and Technology Major Project (No. 2022ZD0115101), National Natural Science Foundation of China Project (No. 623B2086, No. U21A20427), Project (No. WU2022A009) from the Center of Synthetic Biology and Integrated Bioengineering of Westlake University and Integrated Bioengineering of Westlake University and Project (No. WU2023C019) from the Westlake University Industries of the Future Research Funding.
@@ A. Influencing Factors
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes]
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [No] We have considered potential societal impacts and have determined that there are no negative societal impacts associated with our work.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [No] We have used publicly available datasets (DAVIS and BIOSNAP) for testing and comparison purposes. Specific licensing information was not mentioned.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [No]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [No] Our work uses publicly available benchmark data, and no additional consent was required.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [No] The public benchmark data used does not contain personally identifiable information or offensive content.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]